# Universal Discrete-Domain Speech Enhancement

### v1

@@ Abstract
In real-world scenarios, speech signals are inevitably corrupted by various types of interference, making speech enhancement (SE) a critical task for robust speech processing. However, most existing SE methods only handle a limited range of distortions, such as additive noise, reverberation, or band limitation, while the study of SE under multiple simultaneous distortions remains limited. This gap affects the generalization and practical usability of SE methods in real-world environments. To address this gap, this paper proposes a novel Universal Discrete-domain SE model called UDSE. Unlike regression-based SE models that directly predict clean speech waveform or continuous features, UDSE redefines SE as a discrete-domain classification task, instead predicting the clean discrete tokens quantized by the residual vector quantizer (RVQ) of a pre-trained neural speech codec. Specifically, UDSE first extracts global features from the degraded speech. Guided by these global features, the clean token prediction for each VQ follows the rules of RVQ, where the prediction of each VQ relies on the results of the preceding ones. Finally, the predicted clean tokens from all VQs are decoded to reconstruct the clean speech waveform. During training, the UDSE model employs a teacher-forcing strategy, and is optimized with cross-entropy loss. Experimental results confirm that the proposed UDSE model can effectively enhance speech degraded by various conventional and unconventional distortions, e.g., additive noise, reverberation, band limitation, clipping, phase distortion, and compression distortion, as well as their combinations. These results demonstrate the superior universality and practicality of UDSE compared to advanced regression-based SE methods.
@@ I Introduction
Speech signals inevitably suffer from various types of distortion in practical applications [1]. For instance, in outdoor environments, speech is frequently corrupted by different types of additive noise. In transmission scenarios, when speech is transmitted between narrowband communication devices, only the low-frequency components are typically preserved, resulting in bandwidth limitation and the introduction of additional compression noise [2]. These interferences significantly degrade the quality of the speech signal, posing challenges to the understanding of the target speech. Speech enhancement (SE) aims to process degraded speech signals using specific techniques to remove noise, reverberation, and other interfering components, with the goal of improving speech clarity and intelligibility [3]. The enhanced high-quality speech can better serve downstream tasks such as speech communication [4, 5, 6], hearing aid devices, and automatic speech recognition (ASR) [7, 8, 9]. High-quality, efficient, and generalizable SE techniques have become key research objectives for scholars worldwide, carrying significant academic and practical value.
@@
In the past decade, with the development of deep learning, neural network-based SE models have significantly surpassed traditional algorithms in terms of enhanced speech quality and have gradually become the mainstream solutions for SE [10]. At present, most neural SE models operate in the continuous domain, formulating SE as a regression task in which the neural network directly predicts clean speech waveforms or continuous spectral features and employs regression-based loss functions. Regression-based continuous-domain neural SE models [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] can be roughly categorized into waveform-modeling-based and spectrum-modeling-based approaches according to their prediction targets and modeling objectives. Waveform-modeling-based models directly predict clean speech waveforms from degraded ones using a single neural network [11, 12, 13, 14, 15, 16, 17, 18]. For example, SEGAN [12] is an early representative model that employs a fully convolutional network to directly predict clean speech waveforms and introduces a discriminator to enable adversarial training. DEMUCS [17] employs a multi-layer convolutional encoder-decoder architecture with U-Net-style skip connections, along with a sequence modeling module applied to the encoder output, to directly predict the clean speech waveform.
@@
However, direct prediction of clean speech waveforms often leads to low generation efficiency and high computational complexity. In contrast, spectrum-model-based models [19, 20, 21, 22, 23, 24, 25, 26, 27] predict continuous clean spectral features and focus on enhancing speech at the frequency level, thereby avoiding direct manipulation of waveforms. Early spectrum-modeling-based SE models only enhance the amplitude spectrum while ignoring phase degradation, resulting in limited speech quality improvement [21, 22, 23]. To further improve phase enhancement, recent studies have proposed directly enhancing the short-time complex spectrum, thereby enabling the implicit enhancement of both amplitude and phase components. For instance, CMGAN [25] uses a Conformer-based [28] backbone to predict the amplitude spectrum and the real and imaginary parts of the short-time complex spectrum, which are then used to reconstruct the enhanced speech. However, these methods still fall short in achieving explicit and accurate phase enhancement. To address this, in our previous work, we proposed MP-SENet [26, 27], a model that performs parallel and direct enhancement of both the amplitude and phase spectra based on anti-wrapping phase prediction [29, 30], demonstrating impressive results.
@@
Recently, the rapid development of neural speech coding technologies has fostered innovation in the field of SE. A few researchers have begun to explore the use of acoustic discrete tokens generated by neural speech codecs for SE, introducing novel discrete-domain approaches that aim to address classification problems rather than regression problems. In this framework, a neural network predicts the quantized acoustic discrete tokens (i.e., classification categories) of clean speech from degraded speech waveforms or continuous features. The enhanced speech is then reconstructed by decoding these predicted discrete tokens. The loss function is defined as the classification loss between the predicted and the ground truth clean discrete tokens. However, these methods are still in the early stages of development and face numerous challenges. For example, most existing work [31, 32, 33, 34] relies on the capabilities of large language models (LLMs) for token predictions, resulting in excessive model complexity that hinders the practical deployment of SE systems. Genhancer [35] employs a relatively simple DF-Conformer [36] architecture but still relies on auxiliary components, such as ASR tools and self-supervised speech models, to extract text and semantic tokens for assisting acoustic token predictions, resulting in a complex and impractical solution.
@@
In the past one to two years, researchers have started to place greater emphasis on universal SE, aiming to improve the robustness and adaptability of SE models across various distortion types and complex scenarios. For example, the URGENT Challenge [37], which was launched in 2024 and has been held for two consecutive editions, aims to advance universal SE techniques targeting multiple distortion types under challenging real-world conditions. However, current regression-based continuous-domain SE models primarily focus on addressing conventional distortions, such as additive noise, reverberation, and band limitation [38, 39, 40]. These models often struggle to handle unconventional distortion types or scenarios involving multiple simultaneous interferences. In theory, classification-based discrete-domain approaches hold promise for achieving universal SE. This is because many continuous-domain models are designed under specific assumptions [24], such as the weak correlation between noise and the speech signal. Such models tend to perform poorly when facing correlated noise (e.g., compression artifacts), where the noise shares strong characteristics with the underlying speech signal. In contrast, discrete-domain models treat SE as the prediction of acoustic discrete tokens, performing this task consistently regardless of the distortion type. This theoretically makes them more flexible in handling diverse and complex distortions. Nonetheless, current research on discrete-domain methods remains largely focused on conventional additive noise, with limited exploration of their universality across a wider range of distortions.
@@
  • We propose a novel classification-based discrete-domain SE approach, offering a new solution to the SE field while promoting the application of coding techniques within this domain.
  • We propose a novel token prediction paradigm that aligns with the RVQ scheme while also providing an effective solution for other discrete-domain speech generation tasks, such as zero-shot text-to-speech (TTS) with LLMs.
@@
This paper is organized as follows. In Section II, we briefly review the advanced neural speech codecs and their applications in discrte-domain speech generation methods. In Section III, we provide the details of our proposed UDSE model. In Section IV, we present our experimental results. Finally, we give conclusions in Section V.
@@ II-A Neural Speech Coding
As a key speech signal processing technology, speech coding aims to efficiently compress and reconstruct speech signals while preserving perceptual quality. A speech codec serves as the concrete implementation of speech coding, compressing speech signals into discrete tokens that can be used to accurately reconstruct the original speech. A typical speech codec consists of an encoder, a quantizer, and a decoder. Speech codecs can generally be divided into two categories: traditional speech codecs and neural speech codecs. Traditional speech codecs rely on signal processing pipelines and carefully crafted design choices. For example, Opus [41] and EVS [42] integrate techniques such as linear predictive coding (LPC), code-excited linear prediction (CELP), and modified discrete cosine transform (MDCT) to achieve efficient speech coding at various bitrates and sampling rates. However, these codecs often require relatively high bitrates to maintain quality, suffer from noticeable degradation at low bitrates, and are more vulnerable to noise.
@@
With the advancement of deep learning, data-driven end-to-end neural speech codecs have demonstrated significantly superior performance compared to traditional approaches. These neural codecs learn deep representations of speech in an adaptive manner, allowing them to capture essential information and eliminate redundancy even at low bitrates. As a result, they offer higher performance and have emerged as the dominant research direction in the field of speech coding. A central component of neural speech codecs is the learnable quantizer [43, 44, 45, 46]. High quantization quality, reflected by the close alignment between the encoder features and their quantized counterparts, enables the decoder to generate higher-quality reconstructed speech. In early work, researchers proposed a vector quantization (VQ) method that utilizes a learnable codebook, where each feature vector is assigned to the codeword with the minimum Euclidean distance, allowing effective representation learning within a differentiable quantization process [47]. Although this approach addresses the gradient continuity issue in neural network training, it often struggles with convergence and incurs high quantization distortion. To address these limitations, RVQ was introduced and successfully applied in building the neural speech codec SoundStream [43]. RVQ connects multiple VQ modules in a residual manner by iteratively computing and quantizing the residual errors from previous quantization steps. Compared to single-VQ methods, RVQ further reduces overall quantization loss through this residual quantization process.
@@
Building on SoundStream [43], which is the predecessor of neural speech codecs with RVQ, researchers have introduced further improvements in model architecture and training strategies, leading to the development of impressive codecs such as EnCodec [44] and DAC [45]. Recently, we have also proposed improved strategies regarding the coding targets, motivated by the observation that directly coding speech waveforms requires multiple downsampling and upsampling steps, leading to high computational complexity. For example, we introduced APCodec [46], which encodes and decodes the amplitude and phase spectra of speech, and further proposed MDCTCodec [48], which regards the MDCT spectrum of speech as coding target. By shifting the coding target from the waveform level to the spectral level, these approaches significantly reduce computational complexity while maintaining high reconstructed speech quality.
II-B Discrete-Domain Speech Generation
The rapid development of neural speech codecs has brought new solutions to downstream speech generation tasks. This is because neural speech codecs discretize speech signals, enabling powerful models that operate on discrete representations, to be applied to speech generation tasks. Taking TTS as an example, traditional systems often adopt a modular pipeline consisting of text analysis, acoustic feature prediction, and vocoding. While effective, these systems rely heavily on hand-crafted features and are limited in their ability to capture nuanced prosody, speaker characteristics, and emotional cues [49, 50]. The discretization of speech enables the application of LLMs to zero-shot TTS, allowing the generation of speech with the desired text and speaker timbre using only a short prompt speech [51, 52, 53, 54, 55]. In this process, the discrete tokens produced by neural speech codecs serve as a bridge between text and speech. This approach offers several advantages, including the ability to capture speaker characteristics with minimal data, facilitate style and emotion transfer, and leverage the powerful language modeling capabilities of LLMs for expressive speech synthesis.
@@ II-B Discrete-Domain Speech Generation
In addition, recent studies have started to focus on applying neural speech codecs to the SE field, constructing novel discrete-domain SE methods. These methods leverage the discrete tokens produced by neural speech codecs as a bridge between degraded and clean speech, enabling SE within a discrete representation space. For example, Genhancer [35] adopts a DF-Conformer as its core architecture and combines discrete tokens with a self-supervised speech model to perform SE in the discrete domain. GenSE [34] employs a hierarchical modeling framework. It first uses an LLM to refine the semantic tokens extracted from degraded speech, producing enhanced semantic representations. These refined semantic tokens, along with the degraded semantic tokens and acoustic tokens of the degraded speech, are then jointly processed by another LLM to generate enhanced acoustic tokens, which are finally decoded to produce the enhanced speech. However, the application of discrete-domain methods in SE remains in its early stages and faces several challenges that hinder practical deployment. Existing approaches often rely on large-scale models and complex operations, and there has been insufficient exploration of their effectiveness for universal SE.
Figure 1: Overview of the proposed UDSE model.
@@ III-B Global Feature Extraction Module
The reason we use the quantized results to extract global features is that the information bottleneck [56] of VQ helps to initially remove the time-invariant features in the degraded speech, which can reduce the difficulty of SE to some extent. Finally, to fully preserve the original information, the quantization outputs from each VQ along with $\bm{E}^{(y)}$ are fed into the feature processor $\phi_{FP}$. Within $\phi_{FP}$, the input features are concatenated, dimension-reduced, and passed through $B_{G}$ Conformer blocks [28], with residual connections applied to each block to link its input and output. This process generates the global feature $\bm{G}^{(y)}$, i.e,
@@ III-C Token Prediction Module
where $d_{0,*}/\hat{d}_{*,*}^{(x)}\in\{1,2,\dots,M\}$. The token number $L$ is determined by the number of frames in $\bm{G}^{(y)}$ and $M$ is the codebook size of a VQ. The token prediction module consists of $N$ token predictors $\phi_{TP_{1}},\dots,\phi_{TP_{N}}$, each of which has a backbone composed of $B_{T}$ Conformer blocks [28] with residual connections.

### v2

@@ Abstract
In real-world scenarios, speech signals are inevitably corrupted by various types of interference, making speech enhancement (SE) a critical task for robust speech processing. However, most existing SE methods only handle a limited range of distortions, such as additive noise, reverberation, or band limitation, while the study of SE under multiple simultaneous distortions remains limited. This gap affects the generalization and practical usability of SE methods in real-world environments. To address this gap, this paper proposes a novel Universal Discrete-domain SE model called UDSE. UDSE primarily works to enhance speech by predicting clean discrete tokens that are quantized by the residual vector quantizer (RVQ) of a pre-trained neural speech codec and are predicted in accordance with the RVQ rules. Specifically, UDSE first extracts global features from the degraded speech. Guided by these global features, the clean token prediction for each VQ follows the rules of RVQ, where the prediction of each VQ relies on the results of the preceding ones. Finally, the predicted clean tokens from all VQs are decoded to reconstruct the clean speech waveform. During training, the UDSE model employs a teacher-forcing strategy, and is optimized with cross-entropy loss. Experimental results confirm that the proposed UDSE model can effectively enhance speech degraded by various conventional and unconventional distortions, e.g., additive noise, reverberation, band limitation, clipping, phase distortion, and compression distortion, as well as their combinations. These results demonstrate the superior universality and practicality of UDSE compared to advanced regression-based SE methods.
@@ I Introduction
Speech signals inevitably suffer from various types of distortion in practical applications [benesty2008microphone]. For instance, in outdoor environments, speech is frequently corrupted by different types of additive noise. In transmission scenarios, when speech is transmitted between narrowband communication devices, only the low-frequency components are typically preserved, resulting in bandwidth limitation and the introduction of additional compression noise [itu2001712]. These interferences significantly degrade the quality of the speech signal, posing challenges to the understanding of the target speech. Speech enhancement (SE) aims to process degraded speech signals using specific techniques to remove noise, reverberation, and other interfering components, with the goal of improving speech clarity and intelligibility [loizou2007speech]. The enhanced high-quality speech can better serve downstream tasks such as speech communication [rabiner1995impact, mahdi2009advances, ince1992overview], hearing aid devices, and automatic speech recognition (ASR) [wang2019bridging, chen2015integration, delcroix2013speech]. High-quality, efficient, and generalizable SE techniques have become key research objectives for scholars worldwide, carrying significant academic and practical value.
@@
In the past decade, with the development of deep learning, neural network-based SE models have significantly surpassed traditional algorithms in terms of enhanced speech quality and have gradually become the mainstream solutions for SE [wang2018supervised]. At present, most neural SE models operate in the continuous domain, formulating SE as a regression task in which the neural network directly predicts clean speech waveforms or continuous spectral features and employs regression-based loss functions. Regression-based continuous-domain neural SE models [tan2018convolutional, pascual2017segan, rethage2018wavenet, wang2021tstnn, luo2019conv, liu2022voicefixer, defossez2020real, serra2022universal, su2021hifi, ephraim2003speech, xu2014regression, kim2020t, yu2022dual, chu2021speech, abdulatif2024cmgan, lu2023mp, lu2025explicit] can be roughly categorized into waveform-modeling-based and spectrum-modeling-based approaches according to their prediction targets and modeling objectives. Waveform-modeling-based models directly predict clean speech waveforms from degraded ones using a single neural network [tan2018convolutional, pascual2017segan, rethage2018wavenet, wang2021tstnn, luo2019conv, liu2022voicefixer, defossez2020real, serra2022universal]. For example, SEGAN [pascual2017segan] is an early representative model that employs a fully convolutional network to directly predict clean speech waveforms and introduces a discriminator to enable adversarial training. DEMUCS [defossez2020real] employs a multi-layer convolutional encoder-decoder architecture with U-Net-style skip connections, along with a sequence modeling module applied to the encoder output, to directly predict the clean speech waveform.
@@
However, direct prediction of clean speech waveforms often leads to low generation efficiency and high computational complexity. In contrast, spectrum-model-based models [su2021hifi, ephraim2003speech, xu2014regression, kim2020t, yu2022dual, chu2021speech, abdulatif2024cmgan, lu2023mp, lu2025explicit] predict continuous clean spectral features and focus on enhancing speech at the frequency level, thereby avoiding direct manipulation of waveforms. Early spectrum-modeling-based SE models only enhance the amplitude spectrum while ignoring phase degradation, resulting in limited speech quality improvement [xu2014regression, kim2020t, yu2022dual]. To further improve phase enhancement, recent studies have proposed directly enhancing the short-time complex spectrum, thereby enabling the implicit enhancement of both amplitude and phase components. For instance, CMGAN [abdulatif2024cmgan] uses a Conformer-based [gulati2020conformer] backbone to predict the amplitude spectrum and the real and imaginary parts of the short-time complex spectrum, which are then used to reconstruct the enhanced speech. However, these methods still fall short in achieving explicit and accurate phase enhancement. To address this, in our previous work, we proposed MP-SENet [lu2023mp, lu2025explicit], a model that performs parallel and direct enhancement of both the amplitude and phase spectra based on anti-wrapping phase prediction [ai2023neural, ai2024low], demonstrating impressive results.
@@
Recently, the rapid development of neural speech coding technologies has fostered innovation in the field of SE. A few researchers have begun to explore the use of acoustic discrete tokens generated by neural speech codecs for SE, introducing novel discrete-domain approaches that aim to address classification problems rather than regression problems. In this framework, a neural network predicts the quantized acoustic discrete tokens (i.e., classification categories) of clean speech from degraded speech waveforms or continuous features. The enhanced speech is then reconstructed by decoding these predicted discrete tokens. The loss function is defined as the classification loss between the predicted and the ground truth clean discrete tokens. However, these methods are still in the early stages of development and face numerous challenges. For example, most existing work [wang2024speechx, xue2024low, wang2024selm, yao2025gense, kang-etal-2025-llase] relies on the capabilities of large language models (LLMs) for token predictions, resulting in excessive model complexity that hinders the practical deployment of SE systems. Genhancer [yang2024genhancer] employs a relatively simple DF-Conformer [koizumi2021df] architecture but still relies on auxiliary components, such as ASR tools and self-supervised speech models, to extract text and semantic tokens for assisting acoustic token predictions, resulting in a complex and impractical solution.
@@
In the past one to two years, researchers have started to place greater emphasis on universal SE, aiming to improve the robustness and adaptability of SE models across various distortion types and complex scenarios. For example, the URGENT Challenge [zhang2024urgent], which was launched in 2024 and has been held for two consecutive editions, aims to advance universal SE techniques targeting multiple distortion types under challenging real-world conditions. However, current regression-based continuous-domain SE models primarily focus on addressing conventional distortions, such as additive noise, reverberation, and band limitation [kim2023hd, yang2024dm, nair2021cascaded]. These models often struggle to handle unconventional distortion types or scenarios involving multiple simultaneous interferences. This is because many continuous-domain models formulate the task as a regression problem [chu2021speech], aiming to learn the underlying noise patterns and to fit an explicit input–output mapping through neural networks. Such models tend to perform poorly when facing correlated noise (e.g., compression artifacts), where the noise shares strong characteristics with the underlying speech signal and cannot be explicitly modeled or fitted. In contrast, classification-based discrete-domain approaches theoretically hold promise for achieving universal SE. These methods formulate the problem as a classification task that maps the input into a latent discrete space. They leverage the strong generative capability of generative models to learn data distributions and then decode waveforms from the classified tokens. Furthermore, the cross-entropy loss used in the discrete domain penalizes incorrect predictions more severely, which helps the model better distinguish between correct and incorrect outputs. This theoretically makes them more flexible in handling diverse and complex distortions. Nonetheless, current research on discrete-domain methods remains largely focused on conventional additive noise, with limited exploration of their universality across a wider range of distortions.
@@
  • The proposed UDSE model introduces a novel token prediction paradigm that aligns with the RVQ scheme and may also provide inspiration for other discrete-domain speech generation tasks.
@@
This paper is organized as follows. In Section II, we briefly review the advanced neural speech codecs and their applications in discrete-domain speech generation methods. In Section III, we provide the details of our proposed UDSE model. In Section IV, we present our experimental results. Finally, we give conclusions in Section V.
@@ II-A Neural Speech Coding
With the rise of deep learning, end-to-end neural speech codecs now markedly outperform traditional methods. They learn adaptive, compact representations that retain essential information and remove redundancy even at low bitrates, making them the dominant direction in speech coding. A central component of neural speech codecs is the learnable quantizer [zeghidour2021soundstream, defossezhigh, kumar2023high, ai2024apcodec]. High quantization quality, reflected by the close alignment between the encoder features and their quantized counterparts, enables the decoder to generate higher-quality reconstructed speech. In early work, researchers proposed a vector quantization (VQ) method that utilizes a learnable codebook, where each feature vector is assigned to the codeword with the minimum Euclidean distance, allowing effective representation learning within a differentiable quantization process [garbacea2019low]. Although this alleviates gradient-continuity issues, it often converges unstably and yields higher distortion. To address these limitations, RVQ was introduced and successfully applied in building the neural speech codec SoundStream [zeghidour2021soundstream]. RVQ connects multiple VQ modules in a residual manner by iteratively computing and quantizing the residual errors from previous quantization steps. Compared to single-VQ methods, RVQ further reduces overall quantization loss through this residual quantization process.
@@
Building on SoundStream [zeghidour2021soundstream], which is the predecessor of neural speech codecs with RVQ, researchers have introduced further improvements in model architecture and training strategies, leading to the development of impressive codecs such as EnCodec [defossezhigh] and DAC [kumar2023high]. Recently, we have also proposed improved strategies regarding the coding targets, motivated by the observation that directly coding speech waveforms requires multiple downsampling and upsampling steps, leading to high computational complexity. For example, we introduced APCodec [ai2024apcodec], which encodes and decodes the amplitude and phase spectra of speech, and further proposed MDCTCodec [jiang2024mdctcodec], which regards the MDCT spectrum of speech as coding target. By shifting the coding target from the waveform level to the spectral level, these approaches significantly reduce computational complexity while maintaining high reconstructed speech quality.
II-B Universal Speech Enhancement
In real-world scenarios, distortion is highly diverse, and SE models designed for a single distortion type have limited effectiveness in practice. Therefore, exploring universal SE methods is of great importance. Recently, several generative SE approaches based on diffusion models have demonstrated strong performance [lu2022conditional, lemercier2023storm, richter2023speech, universe, universe++]. For example, UniverSE++ [universe++] combines score-based diffusion and adversarial training: the score-based diffusion model uses stochastic differential equations to gradually transform clean speech into noise, while in the reverse process, it learns to recover the clean speech. Adversarial training further helps the model generate more natural-sounding speech, achieving excellent performance across various distortion types.
@@ II-B Universal Speech Enhancement
In addition to the abovementioned diffusion-based SE methods, recent studies have started to focus on applying neural speech codecs to the SE field, constructing novel discrete-domain SE approaches. This type of method has the potential to achieve universal SE by equivalently converting enhancement tasks of any distortion type into a classification problem over discrete representations. The discrete-domain SE leverages the discrete tokens produced by neural speech codecs as a bridge between degraded and clean speech, enabling SE within a discrete representation space. For example, Genhancer [yang2024genhancer] adopts a DF-Conformer as its core architecture and combines discrete tokens with a self-supervised speech model to perform SE in the discrete domain. GenSE [yao2025gense] employs a hierarchical modeling framework. It first uses an LLM to refine the semantic tokens extracted from degraded speech, producing enhanced semantic representations. These refined semantic tokens, along with the degraded semantic tokens and acoustic tokens of the degraded speech, are then jointly processed by another LLM to generate enhanced acoustic tokens, which are finally decoded to produce the enhanced speech. However, existing approaches often depend on large-scale models and complicated processing pipelines, which undermines their practical applicability, and there has also been insufficient exploration of their effectiveness for universal SE.
Figure 1: Overview of the proposed UDSE’s inference process.
@@ III-B Global Feature Extraction Module
Finally, to fully preserve the original information, the quantization outputs from each VQ along with $\bm{E}^{(y)}$ are fed into the feature processor $\phi_{FP}$. Within $\phi_{FP}$, the input features are concatenated, dimension-reduced, and passed through $B_{G}$ Conformer blocks [gulati2020conformer], with residual connections applied to each block to link its input and output. This process generates the global feature $\bm{G}^{(y)}$, i.e,
@@ III-C Token Prediction Module
where $d_{0,*}/\hat{d}_{*,*}^{(x)}\in\{1,2,\dots,M\}$. The token number $L$ is determined by the number of frames in $\bm{G}^{(y)}$ and $M$ is the codebook size of a VQ. The token prediction module consists of $N$ token predictors $\phi_{TP_{1}},\dots,\phi_{TP_{N}}$, each of which has a backbone composed of $B_{T}$ Conformer blocks [gulati2020conformer] with residual connections.